An x-ray photon of initial frequency 3.0x10^-19 sec^-1 collides with an electron and is scattered through 60
1
Expert's answer
2021-09-12T19:01:06-0400
"\lambda'-\lambda=\frac{h}{mc}(1-\cos{\theta})\\
\frac{1}{f'}-\frac{1}{f}=\frac{h}{mc^2}(1-\cos{\theta})
\\
\frac{1}{f'}-\frac{1}{3\cdot10^{-19}}=\frac{6.63\cdot10^{-34}}{(9.1\cdot10^{-31})(3\cdot10^{8})^2}(1-\cos{60})\\f'=3\cdot10^{-19}\ Hz"